OpenStudy (anonymous):

Solve for x: -5(x - 3) = 30 how do I solve this?

OpenStudy (blurbendy):

Start by distributing the -5 -5x + 15 = 30 -5x = 15 x = - 3 Put -3 back into x for the original equation to check your answer.

OpenStudy (anonymous):

thank you so much :)